### v1

@@ 0.C.2 Solver Performance on Individual Benchmarks
Alt. Loops REI.id528.ep89177 SAT 0.8 7.0 SAT 53.8 53.1
Alt. Loops REI.id528.ep89331 SAT 0.8 6.7 SAT 158.9 48.1
Alt. Loops REI.id528.ep90341 SAT 0.7 6.5 SAT 36.3 39.8
Alt. Loops REI.id528.ep95627 SAT 0.7 6.9 SAT 0.1 8.3
Alt. Loops REI.id530.ep82130 SAT 0.4 6.5 SAT 0.2 13.2
Alt. Loops REI.id530.ep82185 SAT 0.7 7.4 SAT 0.3 12.9
Alt. Loops REI.id530.ep82207 SAT 0.5 6.7 SAT 0.2 13.1
Alt. Loops REI.id530.ep91636 SAT 0.6 7.0 SAT 0.4 31.3
Alt. Loops REI.id530.ep94463 SAT 0.5 6.9 SAT 0.2 12.8
Alt. Loops REI.id535.ep76894 UNS 322.8 9.7 UNS 54.5 43.6
Alt. Loops REI.id535.ep76916 UNS 23.1 9.8 UNS 51.4 39.6
Alt. Loops REI.id535.ep87193 SAT 0.8 7.0 SAT 0.2 13.9
Alt. Loops REI.id535.ep87226 SAT 0.9 7.1 SAT 0.5 22.8
Alt. Loops REI.id535.ep91323 SAT 25.2 9.4 UNS 176.3 44.0
Alt. Loops REI.id540.ep57595 UNS 1472.6 10.9 UNK 3601.0 65.2
Alt. Loops REI.id540.ep57610 UNS 1646.9 11.0 UNK 3601.0 66.9
Alt. Loops REI.id540.ep57775 UNS 2759.4 11.2 UNK 3601.0 68.2
Alt. Loops REI.id540.ep85904 SAT 0.9 7.0 SAT 762.1 68.3
Alt. Loops REI.id540.ep94690 SAT 0.8 7.1 SAT 0.1 12.2
Alt. Loops REI.id549.ep85789 SAT 0.7 6.9 SAT 0.2 15.2
Alt. Loops REI.id549.ep87887 SAT 510.8 11.4 SAT 1509.6 64.1
Alt. Loops REI.id549.ep93366 SAT 0.9 6.9 SAT 28.6 51.1
Alt. Loops REI.id549.ep99098 SAT 0.9 6.9 SAT 0.3 16.4
Alt. Loops REI.id549.ep99331 SAT 0.7 6.7 SAT 0.1 8.2
Alt. Loops REI.id67.ep82981 UNS 26.2 11.2 UNS 48.7 33.4
Alt. Loops REI.id67.ep84585 UNS 54.5 10.7 UNS 98.8 39.6
Alt. Loops REI.id67.ep87611 UNS 53.7 10.9 UNS 29.7 34.7
Alt. Loops REI.id67.ep95115 SAT 0.6 6.8 SAT 111.2 41.5
Alt. Loops REI.id67.ep99214 SAT 1.0 7.1 SAT 0.8 35.5
Alt. Loops REI.id68.ep95777 SAT 0.6 7.0 SAT 0.4 20.0
Alt. Loops REI.id68.ep95821 SAT 0.5 6.9 SAT 0.2 10.8
Alt. Loops REI.id68.ep95854 SAT 0.5 6.6 SAT 0.3 16.7
Alt. Loops REI.id68.ep95942 SAT 0.4 6.7 SAT 0.2 11.4
Alt. Loops REI.id68.ep98804 SAT 0.8 7.3 SAT 0.3 23.0
Alt. Loops REI.id73.ep74414 SAT 299.2 11.7 SAT 81.0 52.7
Alt. Loops REI.id73.ep74469 SAT 262.1 11.9 SAT 23.7 50.4
Alt. Loops REI.id73.ep82370 UNS 325.1 11.7 UNS 1166.3 58.1
Alt. Loops REI.id73.ep84023 UNS 296.5 11.2 UNS 946.3 54.9
Alt. Loops REI.id73.ep91504 SAT 0.6 6.8 SAT 153.7 53.7
Alt. Loops REI.id76.ep90543 UNS 288.1 10.6 UNS 2095.2 62.3
Alt. Loops REI.id76.ep91480 UNS 255.0 11.3 UNS 1264.9 62.4
Alt. Loops REI.id76.ep92143 UNS 289.1 10.5 UNS 411.0 46.4
Alt. Loops REI.id76.ep98692 UNS 270.3 12.0 UNS 1011.0 54.6
Alt. Loops REI.id76.ep99198 UNS 342.9 11.7 UNS 1862.4 56.5
Alt. Loops REI.id77.ep49509 UNS 39.8 9.3 UNS 69.4 37.8
Alt. Loops REI.id77.ep75183 UNS 38.7 9.2 UNS 102.2 38.4
Alt. Loops REI.id77.ep91446 UNS 49.6 9.6 UNS 165.0 41.0
Alt. Loops REI.id77.ep99119 UNS 67.1 9.4 UNS 147.6 38.8
Alt. Loops REI.id77.ep99229 UNS 68.9 10.0 UNS 144.3 40.5
Alt. Loops REI.id81.ep89983 UNS 209.4 11.2 UNK 3601.0 59.4
Alt. Loops REI.id81.ep91408 UNS 204.4 10.7 UNK 3601.0 60.9
Alt. Loops REI.id81.ep91571 UNS 220.4 10.8 UNS 2485.5 56.8
Alt. Loops REI.id81.ep97775 UNS 313.1 11.4 UNS 3578.0 61.4
Alt. Loops REI.id81.ep97833 UNS 199.5 11.1 UNK 3601.0 57.0
Alt. Loops REI.id90.ep84292 UNS 2081.5 9.9 UNK 3601.0 60.4
Alt. Loops REI.id90.ep94747 UNS 246.2 11.2 UNK 3601.0 59.0
Alt. Loops REI.id90.ep97680 UNS 194.2 10.2 UNK 3601.0 54.8
Alt. Loops REI.id90.ep97955 UNS 219.2 10.9 UNK 3601.0 59.9
Alt. Loops REI.id90.ep99033 UNS 244.9 11.3 UNK 3601.0 58.0
Alt. Loops REI.id91.ep54663 UNS 43.6 9.6 UNS 1790.1 53.0
Alt. Loops REI.id91.ep58150 UNS 37.9 9.0 UNK 3601.0 55.4
Alt. Loops REI.id91.ep58172 UNS 32.6 8.7 UNS 2933.8 56.0
Alt. Loops REI.id91.ep75275 UNS 133.9 10.1 UNS 649.1 48.4
Alt. Loops REI.id91.ep95105 UNS 172.4 10.1 UNS 1257.0 56.5
DeepCert ind0.target6.eps0.2 UNS 10.4 412.4 UNS 268.9 603.7
DeepCert ind0.target6.eps0.4 UNS 10.3 425.5 UNS 830.0 2053.0
DeepCert ind0.target6.eps0.6 UNS 25.1 582.4 UNS 770.8 1252.2
DeepCert ind14.target6.eps0.2 UNS 10.3 412.9 UNS 457.9 1081.1
DeepCert ind14.target6.eps0.4 SAT 26.9 572.0 SAT 617.3 1421.2
DeepCert ind14.target6.eps0.6 SAT 25.2 580.8 SAT 428.7 1086.1
DeepCert ind15.target6.eps0.2 UNS 10.2 410.0 UNS 398.0 602.4
DeepCert ind15.target6.eps0.4 UNS 38.0 571.3 UNS 476.5 951.0
DeepCert ind15.target6.eps0.6 SAT 24.5 579.3 SAT 177.8 753.4
DeepCert ind19.target6.eps0.2 UNS 10.7 409.2 UNS 442.5 938.9
DeepCert ind19.target6.eps0.4 UNS 30.6 573.9 UNS 532.1 1419.9
DeepCert ind19.target6.eps0.6 SAT 26.4 582.6 SAT 419.0 754.9
DeepCert ind20.target6.eps0.2 UNS 10.6 413.6 UNS 352.1 603.2
DeepCert ind20.target6.eps0.4 UNS 26.6 567.3 UNS 1568.0 2244.6
DeepCert ind20.target6.eps0.6 SAT 24.6 579.8 SAT 436.5 1403.0
DeepCert ind21.target6.eps0.2 UNS 11.0 412.6 UNS 366.8 601.6
DeepCert ind21.target6.eps0.4 UNS 30.2 571.0 UNS 1054.4 1889.6
DeepCert ind21.target6.eps0.6 SAT 27.0 582.3 SAT 874.6 2369.0
DeepCert ind22.target6.eps0.2 UNS 10.6 408.4 UNS 261.3 602.2
DeepCert ind22.target6.eps0.4 UNS 26.0 567.8 UNS 698.6 943.7
DeepCert ind22.target6.eps0.6 SAT 26.5 576.0 SAT 376.1 3149.2
DeepCert ind23.target6.eps0.2 UNS 10.8 412.4 UNS 615.3 1075.1
DeepCert ind23.target6.eps0.4 UNS 32.7 573.7 UNS 727.4 1255.8
DeepCert ind23.target6.eps0.6 SAT 24.9 579.4 SAT 162.7 910.2
DeepCert ind24.target6.eps0.2 UNS 10.5 411.7 UNS 829.5 1259.0
DeepCert ind24.target6.eps0.4 UNS 40.3 569.8 UNS 1477.3 1929.0
DeepCert ind24.target6.eps0.6 SAT 26.2 574.3 SAT 272.1 933.5
DeepCert ind25.target6.eps0.2 UNS 11.0 413.6 UNS 933.3 1572.4
DeepCert ind25.target6.eps0.4 SAT 25.8 571.3 SAT 293.6 925.2
DeepCert ind25.target6.eps0.6 SAT 24.9 580.8 SAT 357.4 1251.2
DeepCert ind26.target6.eps0.2 UNS 10.6 413.0 UNS 515.5 781.7
DeepCert ind26.target6.eps0.4 UNS 25.1 571.1 UNS 1136.0 1887.4
DeepCert ind26.target6.eps0.6 SAT 25.1 579.7 SAT 1453.1 1568.0
DeepCert ind27.target6.eps0.2 UNS 10.3 408.0 UNS 500.2 781.4
DeepCert ind27.target6.eps0.4 SAT 24.5 571.3 SAT 628.1 1271.7
DeepCert ind27.target6.eps0.6 SAT 25.2 579.6 SAT 905.0 2196.7
DeepCert ind28.target6.eps0.2 UNS 10.3 412.3 UNS 714.5 945.7
DeepCert ind28.target6.eps0.4 UNS 30.6 571.4 UNS 433.7 939.1
DeepCert ind28.target6.eps0.6 SAT 24.7 579.8 SAT 437.0 1892.1
DeepCert ind29.target6.eps0.2 UNS 10.7 408.6 UNS 597.4 782.6
DeepCert ind29.target6.eps0.4 SAT 24.9 571.6 SAT 687.2 1112.0
DeepCert ind29.target6.eps0.6 SAT 24.7 580.0 SAT 461.1 1248.9
DeepCert ind31.target6.eps0.2 UNS 25.6 551.3 UNS 549.4 949.7
DeepCert ind31.target6.eps0.4 SAT 24.2 565.1 SAT 309.5 1089.2
DeepCert ind31.target6.eps0.6 SAT 23.9 574.7 SAT 366.0 1268.6
DeepCert ind33.target6.eps0.2 UNS 24.0 550.0 UNS 804.9 1414.6
DeepCert ind33.target6.eps0.4 SAT 24.1 560.0 SAT 211.3 1409.0
DeepCert ind33.target6.eps0.6 SAT 24.0 573.9 SAT 357.8 757.1
DeepCert ind39.target6.eps0.2 UNS 9.9 406.6 UNS 300.2 603.1
DeepCert ind39.target6.eps0.4 UNS 30.6 559.5 UNS 1185.9 3033.1
DeepCert ind39.target6.eps0.6 UNS 33.0 573.1 UNS 1742.9 2882.3
DeepCert ind42.target6.eps0.2 UNS 10.7 406.5 UNS 603.8 940.5
DeepCert ind42.target6.eps0.4 UNS 39.0 563.5 UNS 1143.6 3472.4
DeepCert ind42.target6.eps0.6 UNS 53.1 572.3 UNS 1596.8 2047.7
DeepCert ind43.target6.eps0.2 UNS 10.7 407.9 UNS 340.0 602.7
DeepCert ind43.target6.eps0.4 UNS 29.7 565.4 UNS 1673.5 2244.1
DeepCert ind43.target6.eps0.6 UNS 33.3 574.9 UNS 1622.2 3670.7
DeepCert ind45.target6.eps0.2 UNS 10.6 406.9 UNS 428.6 781.9
DeepCert ind45.target6.eps0.4 UNS 31.6 563.4 UNS 1208.6 2048.1
DeepCert ind45.target6.eps0.6 UNS 25.5 572.8 UNS 2229.4 3658.6
DeepCert ind46.target6.eps0.2 UNS 9.8 404.5 UNS 1287.0 1890.3
DeepCert ind46.target6.eps0.4 UNS 26.2 559.8 UNS 1902.9 2328.9
DeepCert ind46.target6.eps0.6 UNS 40.0 565.2 UNS 951.6 2042.0
DeepCert ind47.target6.eps0.2 UNS 10.0 403.7 UNS 631.0 941.5
DeepCert ind47.target6.eps0.4 UNS 24.9 558.5 UNS 736.4 1258.9
DeepCert ind47.target6.eps0.6 UNS 31.8 568.1 UNS 1700.8 1771.6
DeepCert ind51.target6.eps0.2 UNS 10.2 404.7 UNS 365.8 1102.2
DeepCert ind51.target6.eps0.4 UNS 25.4 557.1 UNS 1807.4 2528.8
DeepCert ind51.target6.eps0.6 UNS 39.5 570.0 UNS 2099.4 1973.8
DeepCert ind53.target6.eps0.2 UNS 10.0 402.6 UNS 376.4 783.7
DeepCert ind53.target6.eps0.4 UNS 24.8 558.0 UNS 1429.3 2523.9
DeepCert ind53.target6.eps0.6 UNS 34.3 566.6 UNS 1480.0 1731.3
DeepCert ind55.target6.eps0.2 UNS 9.5 400.3 UNS 235.3 600.9
DeepCert ind55.target6.eps0.4 UNS 24.7 550.6 UNS 1388.7 1972.8
DeepCert ind55.target6.eps0.6 UNS 30.5 563.8 UNS 875.4 2053.1
DeepCert ind56.target6.eps0.2 UNS 10.1 401.2 UNS 579.3 1101.2
DeepCert ind56.target6.eps0.4 UNS 23.5 550.8 UNS 809.9 1137.7
DeepCert ind56.target6.eps0.6 UNS 38.9 564.6 UNS 1320.3 1728.9
DeepCert ind6.target6.eps0.2 UNS 10.8 410.2 UNS 417.5 1098.1
DeepCert ind6.target6.eps0.4 UNS 24.2 566.0 UNS 701.0 1576.8
DeepCert ind6.target6.eps0.6 SAT 27.4 574.4 SAT 734.7 1255.1
DeepCert ind60.target6.eps0.2 UNS 8.3 357.1 UNS 543.8 604.8
DeepCert ind60.target6.eps0.4 UNS 8.4 368.1 UNS 302.9 599.8
DeepCert ind60.target6.eps0.6 UNS 24.7 503.1 UNS 3366.6 2526.6
DeepCert ind66.target6.eps0.2 UNS 7.4 332.3 UNS 465.4 601.4
DeepCert ind66.target6.eps0.4 SAT 15.0 454.2 SAT 1079.6 1255.9
DeepCert ind66.target6.eps0.6 SAT 16.1 463.5 SAT 1648.3 2406.2
DeepCert ind67.target6.eps0.2 UNS 7.0 333.1 UNS 503.2 602.8
DeepCert ind67.target6.eps0.4 UNS 20.0 457.7 UNS 994.9 1110.3
DeepCert ind67.target6.eps0.6 SAT 17.1 466.9 SAT 2728.2 2375.9
DeepCert ind68.target6.eps0.2 UNS 6.2 304.1 UNS 722.2 607.0
DeepCert ind68.target6.eps0.4 UNS 6.6 313.9 UNS 679.7 949.3
DeepCert ind68.target6.eps0.6 UNS 16.1 423.4 UNS 2760.5 1776.7
DeepCert ind69.target6.eps0.2 UNS 6.3 318.5 UNS 583.5 604.9
DeepCert ind69.target6.eps0.4 UNS 14.5 437.6 UNS 657.7 948.4
DeepCert ind69.target6.eps0.6 UNS 16.6 446.0 UNS 3417.4 2102.0
DeepCert ind7.target6.eps0.2 UNS 10.1 411.1 UNS 394.2 781.0
DeepCert ind7.target6.eps0.4 UNS 26.0 569.1 UNS 1031.4 1734.9
DeepCert ind7.target6.eps0.6 SAT 27.2 573.7 SAT 1410.3 2172.1
DeepCert ind70.target6.eps0.2 UNS 6.4 321.3 UNS 721.9 604.4
DeepCert ind70.target6.eps0.4 UNS 12.9 438.6 UNS 756.1 787.1
DeepCert ind70.target6.eps0.6 UNS 18.3 447.7 UNS 2683.7 1617.1
DeepCert ind71.target6.eps0.2 UNS 6.8 325.3 UNS 731.5 607.8
DeepCert ind71.target6.eps0.4 UNS 7.3 340.8 UNS 1426.9 1421.8
DeepCert ind71.target6.eps0.6 UNS 17.9 463.3 UNS 744.2 941.4
DeepCert ind72.target6.eps0.2 UNS 7.6 336.0 UNS 536.4 604.2
DeepCert ind72.target6.eps0.4 UNS 7.6 351.9 UNS 1011.7 1107.8
DeepCert ind72.target6.eps0.6 UNS 18.1 479.6 UNS 934.7 944.2
[…]

### v2

@@ 0.C.2 Solver Performance on Individual Benchmarks
Alt. Loops REI.id528.ep89177 SAT 0.9 6.5 SAT 53.8 53.1
Alt. Loops REI.id528.ep89331 SAT 0.8 6.9 SAT 158.9 48.1
Alt. Loops REI.id528.ep90341 SAT 1.0 6.6 SAT 36.3 39.8
Alt. Loops REI.id528.ep95627 SAT 0.5 6.3 SAT 0.1 8.3
Alt. Loops REI.id530.ep82130 SAT 0.5 6.4 SAT 0.2 13.2
Alt. Loops REI.id530.ep82185 SAT 0.5 6.7 SAT 0.3 12.9
Alt. Loops REI.id530.ep82207 SAT 0.5 6.4 SAT 0.2 13.1
Alt. Loops REI.id530.ep91636 SAT 0.5 6.8 SAT 0.4 31.3
Alt. Loops REI.id530.ep94463 SAT 0.6 6.7 SAT 0.2 12.8
Alt. Loops REI.id535.ep76894 UNSAT 22.1 9.7 UNSAT 54.5 43.6
Alt. Loops REI.id535.ep76916 UNSAT 58.5 9.3 UNSAT 51.4 39.6
Alt. Loops REI.id535.ep87193 SAT 0.8 6.8 SAT 0.2 13.9
Alt. Loops REI.id535.ep87226 SAT 0.6 6.9 SAT 0.5 22.8
Alt. Loops REI.id535.ep91323 UNSAT 16.7 9.4 UNSAT 176.3 44.0
Alt. Loops REI.id540.ep57595 UNSAT 1415.4 10.5 UNK 3601.0 65.2
Alt. Loops REI.id540.ep57610 UNSAT 1984.2 10.8 UNK 3601.0 66.9
Alt. Loops REI.id540.ep57775 UNSAT 1625.2 10.5 UNK 3601.0 68.2
Alt. Loops REI.id540.ep85904 SAT 0.8 6.8 SAT 762.1 68.3
Alt. Loops REI.id540.ep94690 SAT 0.5 6.2 SAT 0.1 12.2
Alt. Loops REI.id549.ep85789 SAT 0.5 6.6 SAT 0.2 15.2
Alt. Loops REI.id549.ep87887 SAT 506.8 10.8 SAT 1509.6 64.1
Alt. Loops REI.id549.ep93366 SAT 0.4 6.4 SAT 28.6 51.1
Alt. Loops REI.id549.ep99098 SAT 1.0 6.8 SAT 0.3 16.4
Alt. Loops REI.id549.ep99331 SAT 0.6 6.5 SAT 0.1 8.2
Alt. Loops REI.id67.ep82981 UNSAT 27.2 10.2 UNSAT 48.7 33.4
Alt. Loops REI.id67.ep84585 UNSAT 53.2 10.6 UNSAT 98.8 39.6
Alt. Loops REI.id67.ep87611 UNSAT 50.7 11.0 UNSAT 29.7 34.7
Alt. Loops REI.id67.ep95115 SAT 0.6 6.6 SAT 111.2 41.5
Alt. Loops REI.id67.ep99214 SAT 0.6 6.8 SAT 0.8 35.5
Alt. Loops REI.id68.ep95777 SAT 0.5 6.8 SAT 0.4 20.0
Alt. Loops REI.id68.ep95821 SAT 0.6 6.5 SAT 0.2 10.8
Alt. Loops REI.id68.ep95854 SAT 0.6 6.8 SAT 0.3 16.7
Alt. Loops REI.id68.ep95942 SAT 0.5 6.4 SAT 0.2 11.4
Alt. Loops REI.id68.ep98804 SAT 0.5 6.6 SAT 0.3 23.0
Alt. Loops REI.id73.ep74414 SAT 328.7 11.0 SAT 81.0 52.7
Alt. Loops REI.id73.ep74469 SAT 308.8 11.4 SAT 23.7 50.4
Alt. Loops REI.id73.ep82370 UNSAT 221.5 11.0 UNSAT 1166.3 58.1
Alt. Loops REI.id73.ep84023 UNSAT 137.7 10.5 UNSAT 946.3 54.9
Alt. Loops REI.id73.ep91504 SAT 0.4 6.2 SAT 153.7 53.7
Alt. Loops REI.id76.ep90543 UNSAT 216.7 10.3 UNSAT 2095.2 62.3
Alt. Loops REI.id76.ep91480 UNSAT 289.6 10.8 UNSAT 1264.9 62.4
Alt. Loops REI.id76.ep92143 UNSAT 208.5 10.6 UNSAT 411.0 46.4
Alt. Loops REI.id76.ep98692 UNSAT 310.5 11.0 UNSAT 1011.0 54.6
Alt. Loops REI.id76.ep99198 UNSAT 315.0 10.9 UNSAT 1862.4 56.5
Alt. Loops REI.id77.ep49509 UNSAT 25.0 8.8 UNSAT 69.4 37.8
Alt. Loops REI.id77.ep75183 UNSAT 29.6 8.9 UNSAT 102.2 38.4
Alt. Loops REI.id77.ep91446 UNSAT 52.7 9.6 UNSAT 165.0 41.0
Alt. Loops REI.id77.ep99119 UNSAT 79.7 9.7 UNSAT 147.6 38.8
Alt. Loops REI.id77.ep99229 UNSAT 73.5 8.9 UNSAT 144.3 40.5
Alt. Loops REI.id81.ep89983 UNSAT 176.4 10.7 UNK 3601.0 59.4
Alt. Loops REI.id81.ep91408 UNSAT 146.9 10.7 UNK 3601.0 60.9
Alt. Loops REI.id81.ep91571 UNSAT 177.1 10.4 UNSAT 2485.5 56.8
Alt. Loops REI.id81.ep97775 UNSAT 204.1 11.0 UNSAT 3578.0 61.4
Alt. Loops REI.id81.ep97833 UNSAT 186.6 10.6 UNK 3601.0 57.0
Alt. Loops REI.id90.ep84292 UNSAT 318.8 10.2 UNK 3601.0 60.4
Alt. Loops REI.id90.ep94747 UNSAT 285.4 10.6 UNK 3601.0 59.0
Alt. Loops REI.id90.ep97680 UNSAT 174.2 10.3 UNK 3601.0 54.8
Alt. Loops REI.id90.ep97955 UNSAT 167.0 10.2 UNK 3601.0 59.9
Alt. Loops REI.id90.ep99033 UNSAT 202.8 10.3 UNK 3601.0 58.0
Alt. Loops REI.id91.ep54663 UNSAT 46.6 8.7 UNSAT 1790.1 53.0
Alt. Loops REI.id91.ep58150 UNSAT 51.6 10.6 UNK 3601.0 55.4
Alt. Loops REI.id91.ep58172 UNSAT 38.4 9.2 UNSAT 2933.8 56.0
Alt. Loops REI.id91.ep75275 UNSAT 104.4 10.7 UNSAT 649.1 48.4
Alt. Loops REI.id91.ep95105 UNSAT 125.5 10.0 UNSAT 1257.0 56.5
DeepCert ind0.target6.eps0.2 UNSAT 8.7 399.7 UNSAT 268.9 603.7
DeepCert ind0.target6.eps0.4 UNSAT 8.5 412.4 UNSAT 830.0 2053.0
DeepCert ind0.target6.eps0.6 UNSAT 24.8 569.3 UNSAT 770.8 1252.2
DeepCert ind14.target6.eps0.2 UNSAT 8.5 402.8 UNSAT 457.9 1081.1
DeepCert ind14.target6.eps0.4 SAT 26.9 561.2 SAT 617.3 1421.2
DeepCert ind14.target6.eps0.6 SAT 24.3 569.9 SAT 428.7 1086.1
DeepCert ind15.target6.eps0.2 UNSAT 8.5 401.4 UNSAT 398.0 602.4
DeepCert ind15.target6.eps0.4 UNSAT 39.4 560.2 UNSAT 476.5 951.0
DeepCert ind15.target6.eps0.6 SAT 24.4 568.7 SAT 177.8 753.4
DeepCert ind19.target6.eps0.2 UNSAT 8.5 402.8 UNSAT 442.5 938.9
DeepCert ind19.target6.eps0.4 UNSAT 33.0 560.7 UNSAT 532.1 1419.9
DeepCert ind19.target6.eps0.6 SAT 25.5 569.5 SAT 419.0 754.9
DeepCert ind20.target6.eps0.2 UNSAT 8.6 403.3 UNSAT 352.1 603.2
DeepCert ind20.target6.eps0.4 UNSAT 25.8 560.9 UNSAT 1568.0 2244.6
DeepCert ind20.target6.eps0.6 SAT 24.3 570.9 SAT 436.5 1403.0
DeepCert ind21.target6.eps0.2 UNSAT 8.7 402.4 UNSAT 366.8 601.6
DeepCert ind21.target6.eps0.4 UNSAT 32.9 560.4 UNSAT 1054.4 1889.6
DeepCert ind21.target6.eps0.6 SAT 27.9 568.9 SAT 874.6 2369.0
DeepCert ind22.target6.eps0.2 UNSAT 8.6 403.0 UNSAT 261.3 602.2
DeepCert ind22.target6.eps0.4 UNSAT 26.5 561.2 UNSAT 698.6 943.7
DeepCert ind22.target6.eps0.6 SAT 27.3 569.6 SAT 376.1 3149.2
DeepCert ind23.target6.eps0.2 UNSAT 8.8 401.9 UNSAT 615.3 1075.1
DeepCert ind23.target6.eps0.4 UNSAT 33.7 560.4 UNSAT 727.4 1255.8
DeepCert ind23.target6.eps0.6 SAT 25.8 568.0 SAT 162.7 910.2
DeepCert ind24.target6.eps0.2 UNSAT 10.0 401.6 UNSAT 829.5 1259.0
DeepCert ind24.target6.eps0.4 UNSAT 35.1 559.1 UNSAT 1477.3 1929.0
DeepCert ind24.target6.eps0.6 SAT 26.8 567.9 SAT 272.1 933.5
DeepCert ind25.target6.eps0.2 UNSAT 8.5 403.3 UNSAT 933.3 1572.4
DeepCert ind25.target6.eps0.4 SAT 24.6 560.6 SAT 293.6 925.2
DeepCert ind25.target6.eps0.6 SAT 23.7 569.4 SAT 357.4 1251.2
DeepCert ind26.target6.eps0.2 UNSAT 8.6 404.0 UNSAT 515.5 781.7
DeepCert ind26.target6.eps0.4 UNSAT 25.1 561.0 UNSAT 1136.0 1887.4
DeepCert ind26.target6.eps0.6 SAT 24.0 568.7 SAT 1453.1 1568.0
DeepCert ind27.target6.eps0.2 UNSAT 8.6 402.4 UNSAT 500.2 781.4
DeepCert ind27.target6.eps0.4 SAT 24.2 561.1 SAT 628.1 1271.7
DeepCert ind27.target6.eps0.6 SAT 24.1 568.9 SAT 905.0 2196.7
DeepCert ind28.target6.eps0.2 UNSAT 8.5 402.4 UNSAT 714.5 945.7
DeepCert ind28.target6.eps0.4 UNSAT 30.7 561.5 UNSAT 433.7 939.1
DeepCert ind28.target6.eps0.6 SAT 24.8 569.2 SAT 437.0 1892.1
DeepCert ind29.target6.eps0.2 UNSAT 8.4 402.1 UNSAT 597.4 782.6
DeepCert ind29.target6.eps0.4 SAT 23.6 561.2 SAT 687.2 1112.0
DeepCert ind29.target6.eps0.6 SAT 25.2 569.3 SAT 461.1 1248.9
DeepCert ind31.target6.eps0.2 UNSAT 26.4 540.6 UNSAT 549.4 949.7
DeepCert ind31.target6.eps0.4 SAT 23.6 554.4 SAT 309.5 1089.2
DeepCert ind31.target6.eps0.6 SAT 24.1 564.2 SAT 366.0 1268.6
DeepCert ind33.target6.eps0.2 UNSAT 20.9 539.6 UNSAT 804.9 1414.6
DeepCert ind33.target6.eps0.4 SAT 23.1 553.2 SAT 211.3 1409.0
DeepCert ind33.target6.eps0.6 SAT 23.8 564.2 SAT 357.8 757.1
DeepCert ind39.target6.eps0.2 UNSAT 8.3 395.9 UNSAT 300.2 603.1
DeepCert ind39.target6.eps0.4 UNSAT 34.5 553.4 UNSAT 1185.9 3033.1
DeepCert ind39.target6.eps0.6 UNSAT 32.7 562.7 UNSAT 1742.9 2882.3
DeepCert ind42.target6.eps0.2 UNSAT 8.4 396.2 UNSAT 603.8 940.5
DeepCert ind42.target6.eps0.4 UNSAT 37.4 552.2 UNSAT 1143.6 3472.4
DeepCert ind42.target6.eps0.6 UNSAT 56.1 561.8 UNSAT 1596.8 2047.7
DeepCert ind43.target6.eps0.2 UNSAT 8.4 398.1 UNSAT 340.0 602.7
DeepCert ind43.target6.eps0.4 UNSAT 29.6 555.1 UNSAT 1673.5 2244.1
DeepCert ind43.target6.eps0.6 UNSAT 31.8 564.1 UNSAT 1622.2 3670.7
DeepCert ind45.target6.eps0.2 UNSAT 8.4 397.3 UNSAT 428.6 781.9
DeepCert ind45.target6.eps0.4 UNSAT 35.5 552.3 UNSAT 1208.6 2048.1
DeepCert ind45.target6.eps0.6 UNSAT 25.6 562.3 UNSAT 2229.4 3658.6
DeepCert ind46.target6.eps0.2 UNSAT 8.1 394.3 UNSAT 1287.0 1890.3
DeepCert ind46.target6.eps0.4 UNSAT 24.6 548.8 UNSAT 1902.9 2328.9
DeepCert ind46.target6.eps0.6 UNSAT 38.4 559.0 UNSAT 951.6 2042.0
DeepCert ind47.target6.eps0.2 UNSAT 8.1 393.1 UNSAT 631.0 941.5
DeepCert ind47.target6.eps0.4 UNSAT 24.4 547.9 UNSAT 736.4 1258.9
DeepCert ind47.target6.eps0.6 UNSAT 30.7 557.4 UNSAT 1700.8 1771.6
DeepCert ind51.target6.eps0.2 UNSAT 8.2 394.8 UNSAT 365.8 1102.2
DeepCert ind51.target6.eps0.4 UNSAT 24.1 551.1 UNSAT 1807.4 2528.8
DeepCert ind51.target6.eps0.6 UNSAT 40.6 559.9 UNSAT 2099.4 1973.8
DeepCert ind53.target6.eps0.2 UNSAT 8.0 392.8 UNSAT 376.4 783.7
DeepCert ind53.target6.eps0.4 UNSAT 24.5 547.4 UNSAT 1429.3 2523.9
DeepCert ind53.target6.eps0.6 UNSAT 34.4 556.1 UNSAT 1480.0 1731.3
DeepCert ind55.target6.eps0.2 UNSAT 7.9 389.9 UNSAT 235.3 600.9
DeepCert ind55.target6.eps0.4 UNSAT 24.4 544.1 UNSAT 1388.7 1972.8
DeepCert ind55.target6.eps0.6 UNSAT 30.8 553.0 UNSAT 875.4 2053.1
DeepCert ind56.target6.eps0.2 UNSAT 8.0 391.0 UNSAT 579.3 1101.2
DeepCert ind56.target6.eps0.4 UNSAT 23.5 544.3 UNSAT 809.9 1137.7
DeepCert ind56.target6.eps0.6 UNSAT 38.8 553.4 UNSAT 1320.3 1728.9
DeepCert ind6.target6.eps0.2 UNSAT 8.7 400.1 UNSAT 417.5 1098.1
DeepCert ind6.target6.eps0.4 UNSAT 23.9 555.5 UNSAT 701.0 1576.8
DeepCert ind6.target6.eps0.6 SAT 26.1 564.1 SAT 734.7 1255.1
DeepCert ind60.target6.eps0.2 UNSAT 6.2 347.4 UNSAT 543.8 604.8
DeepCert ind60.target6.eps0.4 UNSAT 6.0 358.8 UNSAT 302.9 599.8
DeepCert ind60.target6.eps0.6 UNSAT 24.4 492.7 UNSAT 3366.6 2526.6
DeepCert ind66.target6.eps0.2 UNSAT 5.1 321.8 UNSAT 465.4 601.4
DeepCert ind66.target6.eps0.4 SAT 14.5 443.9 SAT 1079.6 1255.9
DeepCert ind66.target6.eps0.6 SAT 14.8 453.3 SAT 1648.3 2406.2
DeepCert ind67.target6.eps0.2 UNSAT 5.2 323.0 UNSAT 503.2 602.8
DeepCert ind67.target6.eps0.4 UNSAT 18.6 447.2 UNSAT 994.9 1110.3
DeepCert ind67.target6.eps0.6 SAT 16.0 456.6 SAT 2728.2 2375.9
DeepCert ind68.target6.eps0.2 UNSAT 4.1 325.0 UNSAT 722.2 607.0
DeepCert ind68.target6.eps0.4 UNSAT 4.4 303.8 UNSAT 679.7 949.3
DeepCert ind68.target6.eps0.6 UNSAT 14.4 413.6 UNSAT 2760.5 1776.7
DeepCert ind69.target6.eps0.2 UNSAT 4.9 308.4 UNSAT 583.5 604.9
DeepCert ind69.target6.eps0.4 UNSAT 12.6 427.8 UNSAT 657.7 948.4
DeepCert ind69.target6.eps0.6 UNSAT 15.7 435.7 UNSAT 3417.4 2102.0
DeepCert ind7.target6.eps0.2 UNSAT 9.5 400.7 UNSAT 394.2 781.0
DeepCert ind7.target6.eps0.4 UNSAT 26.4 558.2 UNSAT 1031.4 1734.9
DeepCert ind7.target6.eps0.6 SAT 26.6 567.1 SAT 1410.3 2172.1
DeepCert ind70.target6.eps0.2 UNSAT 5.1 310.3 UNSAT 721.9 604.4
DeepCert ind70.target6.eps0.4 UNSAT 11.9 428.5 UNSAT 756.1 787.1
DeepCert ind70.target6.eps0.6 UNSAT 17.1 437.9 UNSAT 2683.7 1617.1
[…]